# Argumentative Experience: Reducing Confirmation Bias on Controversial Issues through LLM-Generated Multi-Persona Debates

### v2

@@ 5.1. RQ1: User Experience in Conceptualizing Controversial Topics through Multi-Persona Debate
From the behavioral data collected in the debate system (as described in Section  4.4), users generated an average of 1.865 debate rounds per topic. For each round, an average of 1.734 personas supporting the topic and 1.972 personas opposing the topic were selected. Users made minimal edits (0.187 personas), clicked thumbs up/down 2.628 times to bookmark content, and accessed the bookmark tab only 0.121 times on average. Full descriptive data are provided in Table  6. As we recognized notable variations in the behavior data among participants, we performed a clustering analysis to better understand users’ engagement patterns and their relationship to both the system design and the topics explored.
@@ 5.1.1. Active observers (Group 0)
Participants in this group frequently used thumbs up / down features to express their opinions while observing the debate, and often reviewed the bookmarked content. They typically engaged in more than two debate rounds and showed moderate involvement in selecting and editing personas, balancing supportive and opposing viewpoints. For example, during the debate topic of “Should humans colonize space?”, P14 remarked “I think there’s a lot of good points made. Mainly it comes down to the tangibility of it and the fact that it kind of does feel like we’d be abandoning the mess that we made here [the Earth].” Because P14 found it feasible for space colonization, this participant often gave thumbs-up to compelling arguments and thumbs-down to weaker ones. Similarly, others in this group proactively express their opinions using the thumbs features and spent more time on the bookmarked tabs. We see this pattern reflected in our envisioned use case (Section 3.1). It is marked by frequent use of feedback features, revisiting bookmarked content, and a balanced consideration of both supportive and opposing perspectives.
5.1.2. Multi-round explorers (Group 1)
@@
Participants in this group were inclined to generate more rounds of debate (up to five) but spend less time reading and digesting arguments presented. Notably, we observed a preference for selecting personas that agreed with the given topic. From qualitative feedback, it appears that this may stem from the uncertainty about the topic. By initially choosing supportive personas, participants began to familiarize themselves with the subject matter. For example, in the debate over “Should humans colonize space?” P17 admitted a lack of knowledge in that area. The participant first selected two supportive personas, an aerospace engineer and a resource economist, to gain insights into the potential benefits of space colonization. As the debate progressed, P17 gradually included personas with different perspectives. This pattern reflects a strategy of building understanding through repetition and confirmation, with a gradual move toward considering alternative viewpoints as participant knowledge grows throughout the debate.
5.1.3. Critical readers (Group 2)
@@
Participants in this group exhibited the highest level of engagement with the system. Compared to other two groups, they experienced the longest debate durations, selected more new personas, and frequently edited these personas. A notable characteristic was to actively seek out personas that disagreed with the topic. These behaviors suggest significant mental effort invested in generating, evaluating, refining personas, and reading content to gain a comprehensive understanding of the topic. This behavior might stem from participant propensity for critical thinking, as indicated from their qualitative interviews. For example, in the debate topic of “Should abortion be legal?” P14 initially supported the legalization of abortion, felt that they hadn’t explored the topic deeply enough “[Although] I’m kind of familiar, but I haven’t really delved in deep to all the different sides so I don’t want to take a side on it [quickly].” Despite holding a pro stance, this participant sought out opposite opinions to gain a more comprehensive view of the controversial issue. This pattern exemplifies how some participants, even with initial stances, actively engage with conflicting perspectives to ensure a balanced understanding of the topic.
@@ 5.2.1. Effects of bias mitigation across systems and human factors
Table 3. Comparison of Interface Measurements for Attitude-consistent and Attitude-consistent Conditions. The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05.
@@
The ANOVA analysis conducted above on content-belief consistency but separately with topic familiarity and confirmation bias tendency reveal a consistent result: there is a clear difference on whether participants reading content consistent or inconsistent with their attitude when using the debate system, as reflected in their eye movements, but no difference found in the baseline. While topic familiarity yields no difference in both systems, confirmation bias tendency resulted in significant difference on the baseline but not the debate system. In combined with the system comparison results shown in the paired t-test (Table 3), we can conclude that the debate system effectively reduces selective exposure and increases user engagement with attitude-inconsistent content, while the baseline system activates confirmation bias tendencies, as evidenced by distinct eye movement patterns.
@@ 5.2.2. Effects of bias mitigation across user behavior patterns
Figure 5. Comparison of normalized fixation duration and fixation counts on attitude-consistent and inconsistent information across user behavior groups.
@@
These results suggest that there is significant difference in the eye movements between behavior clusters within the debate system, As content-belief consistency did not significantly influence fixation counts, this suggests that user engagement levels were more closely tied to their behavior patterns when using the debate system rather than whether participants would read content associated with their prior beliefs. Additionally, as post-hoc analysis revealed that Critical Readers (group 2) had significantly higher fixation counts compared to the other groups, this suggests that Critical Readers were more visually engaged with the content. By associating with qualitative results, describing how Critical Readers actively engaged with various personas and generated content (Section  5.1), we can conclude that certain behavioral pattern, i.e., Critical Readers, show a clear sign of engagement demonstrated in their eye movement data.
@@ 5.3. RQ3: Comparison of User Belief Changes between Retrieval-based and Multi-persona Debate System
In this section, we first examine whether participant belief changed or remained unchanged across two systems and the influence of topic familiarity and confirmation bias tendency (Section 5.3.1, similar to our approach for RQ2). Afterward, we explore in greater depth in terms of bidirectional belief changes, i.e., whether participants became more or less supportive, because this helps suggest whether there is a shift toward increased or decreased polarization (Section 5.3.2). As the belief change data does not follow a normal distribution, we applies non-parametric test for analysis.
@@ 5.3.1. Effects of non-directional belief changes across systems and human factors
By visually presenting participant belief changes based on their topic familiarity (Figure 6, the two plots on the left) and confirmation bias tendency (Figure 6, the two plots on the right), we observed a clear trend indicating that participants with higher familiarity with the topic were less likely to change their beliefs. However, there is a less obvious trend shown in confirmation bias tendency. We conducted a generalized linear mixed-effects model analysis to quantify the statistical effects (model details were presented in Appendix F.2). The overall result demonstrates that there is a statistically significant difference across the various levels of topic familiarity and confirmation bias tendency on whether participants changed belief or remained unchanged, while the difference is not shown in interface conditions. We briefly describe the mixed effects for each variable in details as follows.
@@
Confirmation bias tendency Confirmation bias tendency shows a statistically significant positive effect on belief change with an estimated coefficient of $0.4338$ ($SE=0.1873$, $z=2.315$ , $p<0.05$). This means that participants with higher levels of confirmation bias tendency are more likely change their beliefs, which is counterintuitive to our understanding. By examining the different proportions of belief changes across CB groups, we found that the low tendency group had the highest individual percentage of belief change (25.6%), but the combined belief change percentages of the medium and high tendency groups (19.1% + 20.6% = 39.7%) were greater than that. This cumulative effect may indicate that participant belief changes might not operate neatly into a three-level continuum of confirmation bias tendency.
Correlation among three variables: As parts of the generalized fixed-effect results, there is no moderate correlation among group interface, topic familiarity, and confirmation bias tendency, reflecting that these variables operated relatively independently in influencing belief change.
@@ 5.3.2. Effects of directional belief changes across systems and human factors
We examined how system conditions influence participant bi-directional belief changes, weighted by their confidence level, and explored this variance on topic familiarity and confirmation bias tendency.
@@
Table 4. Bi-directional belief change between baseline and experiment condition: a larger number of participants became less supportive after using the experiment system compared to the baseline. The weighted change for the experiment was significantly lower than that of the baseline, suggesting that the experiment is more likely to lead participants to become less supportive of their initial stance. The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05
Figure 7. Bidirectional belief changes box plots for the baseline and experimental interfaces across topic familiarity and confirmation biases tendency: Over both factors, while many participants did not alter their beliefs, which reduced the overall magnitude of changes in the box plots, for participants who did change their beliefs, the multi-persona debate system was more likely to lead participants to become less supportive than the baseline.
System comparison: By mapping bi-directional belief changes — less supportive, more supportive, and unchanged — on both systems (Table 4), we found that although many participants did not change their beliefs ($Retrieval_{unchanged}$ = $68.2\%$ , $Debate_{unchanged}$ = $62.5\%$), more participants became less supportive after using the debate than the baseline ($Debate_{less}$ = $24.4\%$ $Retrieval_{less}$ = $11.8\%$). Additionally, by conducting Wilcoxon sign-ranked test on the weighted belief change (in Table 4, the last row, where the mean value ¡0 indicates becoming less supportive and ¿0 indicates becoming more supportive), the weighted belief change for the debate was significantly lower than that of the baseline ($Debate_{weightedchange}$ = $-0.34$ $Retrieval_{weightedchange}$ = $0.30$, $X^{2}=1040.5$, $p<0.05$).
Topic familiarity: By visualizing participant weighted score for belief change across different levels of topic familiarity (Figure 7, the two plots on the left), we found many participants did not alter their beliefs, which reduced the overall magnitude of changes shown in the box plots. However, among participants who did change their beliefs, there is a lowered central tendency (median and the interquartile range) in the experiment compared to the baseline. This expands on earlier findings of significant difference across levels of topic familiarity (Section 5.3.1), demonstrating that, for participants who finally changed their belief, the debate system led them to become less supportive across different familiarity levels. Additionally, we also found the weighted belief change vary depending on different topics (additional analysis and findings are presented in Appendix F.3).
@@
By combining the statistical results in both Section 5.3.1 and 5.3.2, we learned that while the system conditions did not yield a statistically significant difference on participant non-directional belief changes (i.e., whether participant belief changed or remained unchanged), for directional belief changes, there was a clear distinction. The debate system was more likely to cause participants to become less supportive of their initial stance, whereas the baseline tended to reinforce prior beliefs, making them more supportive. Additionally, user belief changes exhibited a statistically significant increase as topic familiarity decreased. While a statistically significant difference was also observed with confirmation bias tendencies, the trend did not fully align with our hypothesis. This discrepancy may be due to the three-level classification of bias tendency scores, indicating a need for further investigation in future studies.
@@ 6. Discussion
6.1. Enhancing Interaction
While our LLM debate system supports several forms of user interaction (e.g., topic selection, persona revision or creation, argument bookmarking, and controlling debate duration), participants offered several ideas to further increase interaction and engagement. Specifically, they sought greater ability to actively participate in and influence the debate.
For example, some participants suggested use of the search bar to allow them to inject their own opinions into an ongoing debate (P2-3). Similarly, some participants misunderstood the thumbs up and down bookmarking icons as allowing them to influence the ongoing debate and expressed a desire for this feature when realizing it was not supported (P26). Others requested the ability to highlight specific quotes from a persona content and prompt the persona to provide further information or supporting evidence (P25-26, 28, 39). A fact-checking feature was suggested to show the most recent evidence from the web for user-highlighted content, aiding in verification of arguments (P39).

### v3

@@ 5.1. RQ1: User Experience in Conceptualizing Controversial Topics through Multi-Persona Debate
In this section, we present two categories of findings: (1) a mixed-methods result combining quantitative and qualitative data to reveal participants’ behavioral patterns (Section 5.1.1) and (2) qualitative themes exploring participants’ motivations for editing personas (Section 5.1.2) and their perceptions of the generation quality (Section 5.1.3). For brevity, we refer to the LLM-based multi-persona debate system simply as “the debate system” in the following results. Similarly, the baseline retrieval-based system is simplified as “the baseline system”.
5.1.1. Clustered behavioral patterns
From the behavioral data collected in the debate system (as described in Section  4.4), users generated an average of 1.865 debate rounds per topic. For each round, an average of 1.734 personas supporting the topic and 1.972 personas opposing the topic were selected. Users made minimal edits (0.187 personas), clicked thumbs up/down 2.628 times to bookmark content, and accessed the bookmark tab only 0.121 times on average. Participants spent an average of 4.82 minutes interacting with and reading the debate content. Full descriptive data are provided in Table  6. As we recognized notable variations in the behavior data among participants, we performed a clustering analysis to better understand users’ engagement patterns and their relationship to both the system design and the topics explored.
@@ 5.1.1. Clustered behavioral patterns
Active observers (Group 0) Participants in this group frequently used thumbs up / down features to express their opinions while observing the debate, and often reviewed the bookmarked content. They typically engaged in more than two debate rounds and showed moderate involvement in selecting and editing personas, balancing supportive and opposing viewpoints. For example, during the debate topic of “Should humans colonize space?”, P14 remarked “I think there’s a lot of good points made. Mainly it comes down to the tangibility of it and the fact that it kind of does feel like we’d be abandoning the mess that we made here [the Earth].” Because P14 found it feasible for space colonization, this participant often gave thumbs-up to compelling arguments and thumbs-down to weaker ones. Similarly, others in this group proactively express their opinions using the thumbs features and spent more time on the bookmarked tabs. We see this pattern reflected in our envisioned use case (Section 3.1). It is marked by frequent use of feedback features, revisiting bookmarked content, and a balanced consideration of both supportive and opposing perspectives.
@@
Multi-round explorers (Group 1) Participants in this group were inclined to generate more rounds of debate (up to five) but spend less time reading and digesting arguments presented. Notably, we observed a preference for selecting personas that agreed with the given topic. From qualitative feedback, it appears that this may stem from the uncertainty about the topic. By initially choosing supportive personas, participants began to familiarize themselves with the subject matter. For example, in the debate over “Should humans colonize space?” P17 admitted a lack of knowledge in that area. The participant first selected two supportive personas, an aerospace engineer and a resource economist, to gain insights into the potential benefits of space colonization. As the debate progressed, P17 gradually included personas with different perspectives. This pattern reflects a strategy of building understanding through repetition and confirmation, with a gradual move toward considering alternative viewpoints as participant knowledge grows throughout the debate.
@@
Critical readers (Group 2) Participants in this group exhibited the highest level of engagement with the system. Compared to other two groups, they experienced the longest debate durations, selected more new personas, and frequently edited these personas. A notable characteristic was to actively seek out personas that disagreed with the topic. These behaviors suggest significant mental effort invested in generating, evaluating, refining personas, and reading content to gain a comprehensive understanding of the topic. This behavior might stem from participant propensity for critical thinking, as indicated from their qualitative interviews. For example, in the debate topic of “Should abortion be legal?” P14 initially supported the legalization of abortion, felt that they hadn’t explored the topic deeply enough “[Although] I’m kind of familiar, but I haven’t really delved in deep to all the different sides so I don’t want to take a side on it [quickly].” Despite holding a pro stance, this participant sought out opposite opinions to gain a more comprehensive view of the controversial issue. This pattern exemplifies how some participants, even with initial stances, actively engage with conflicting perspectives to ensure a balanced understanding of the topic.
@@
5.1.2. Motivations for persona edits
From the post-task interviews, we found three additional factors that participants frequently mentioned in selecting and revising personas, including persona expertise, individuality, and the overall diversity of personas when selecting and editing them to construct the debate. These factors were closely tied to how participants conceptualized different controversial topics being discussed.
First, many participants preferred selecting both expert and individual personas. Expert personas were more trusted to provide factual information, evidence, and other logical and reliable arguments (P11, 14-15, 22, 24, 33, 36, 38-39). P39 explained that for the topic of “TikTok ban,” experts could provide opinions rather than emotional responses “Parents who are concerned about their kid is great, but I don’t expect them to have all of the data necessary to make an informed decision. I thought there would be experts discussing things logically and providing evidence-based information, as opposed to information driven by emotions.” In contrast, individual personas could provide direct and authentic arguments (P11, 22, 24-25, 33, 38). For example, P38 said, “Whether TikTok should be banned from school. I really want people who are in the school specifically because it’s important to see the people who are actually being affected to know their pain points. So that’s why I picked the teacher, the student, and the administrator.
Additionally, as participants were aware that potential bias could underlie different personas, some aimed to create a balanced representation of opinions by selecting personas that supported different sides of the topic. On the one hand, this approach could be an effective strategy for generating counterarguments (P14, 24, 26, 28, 33, 35-36), as P36 explained, “I tried to balance the number of people who supported the topic with those who opposed it, to sort of weigh which perspective truly aligns with my own thinking and against it.” On the other hand, some would deliberately balance opinions that might be disproportionately represented on one side in real-world discussions (P22, 24, 33). P33 pointed out that “If I’ve been hearing a lot of one point of view, I’ll try to keep generating a persona until I find one that opposes it.
5.1.3. Perceptions of generation quality
Participants employed two main strategies to evaluate the quality of the generated content. First, they assessed whether the persona arguments, opinions, and use cases aligned logically with their customized prompts (e.g., political alignment, expertise). Second, they examined the concreteness, reasonableness, and creativity of the content. Many of them emphasized that this dual-level evaluation helped them further assess the reliability of the multi-persona debate system, particularly given their awareness of potential LLM hallucinations.
For example, participants prioritized arguments supported by contextual background knowledge. When content lacked logical coherence, they often dismissed it as unconvincing and retained their original stance (P19, 29, 30). P30 explained that “I was really curious to see really compelling arguments. When I saw they mostly centered on a trickle-down economy [Topic: ‘Corporate income tax raise’], which I think there’s a lot of evidence against, I didn’t find them strong enough to change my opinion on the topic.
Repetitive or irrelevant debate content further diminished perceived usefulness. For example, some participants criticized exchanges that failed to address the core topic across multiple rounds (P13, 17, 21, 26, 35-36, 39-40). This issue extended to the baseline retrieval system, where irrelevant sourced materials, even when citations were provided, make them feel less useful (P13, 22, 24, 27-28, 39).
Conversely, participants perceived the system as highly useful when debates met three conditions: (1) the content fostered deeper engagement with the topic (P14, 28); (2) it introduced novel, perspective-shifting knowledge (P11–12, 14–15, 28, 30); or (3) it provoked critical self-reflection through creative or unexpected arguments (P14, 24, 34). For example, P30 revised the stance on animal testing after encountering arguments highlighting non-animal technological alternatives, noting: “The argument that we have technology capable of conducting similar tests and obtaining comparable data without involving or harming animals was something I hadn’t considered initially, and it shifted my stance.” Similarly, P24 reconsidered their opposition to violent video games when debates emphasized their potential benefits, “we aren’t inherently violent, and the inability to express aggression doesn’t mean we shouldn’t play.
@@ 5.2.1. Effects of bias mitigation across systems and human factors
Table 3. Comparison of total fixation counts and fixation durations in the interfaces for Attitude-consistent and Attitude-consistent Conditions. The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05.
@@
[…]
Comparison of total fixation duration on attitude-consistent and inconsistent information
@@ 5.2.2. Effects of bias mitigation across user behavior patterns
Figure 5. Comparison of normalized fixation duration and fixation counts on attitude-consistent and inconsistent information across user behavior groups while using the debate system.
@@
Comparison of normalized fixation duration and fixation counts on attitude-consistent and inconsistent information across user behavior groups while using the debate system
[…]
@@ 5.3. RQ3: Comparison of User Belief Changes between Retrieval-based and Multi-persona Debate System
In this section, we first examine whether participant belief changed or remained unchanged across two systems and the influence of topic familiarity and confirmation bias tendency (Section 5.3.1, similar to our approach for RQ2). Afterward, we explore in greater depth in terms of directional belief changes, i.e., whether participants became more or less supportive (Section 5.3.2). As the belief change data does not follow a normal distribution, we apply a non-parametric test for analysis.
@@ 5.3.1. Effects of non-directional belief changes across systems and human factors
By visually presenting participant belief changes based on their topic familiarity (Figure 6, the two plots on the left) and confirmation bias tendency (Figure 6, the two plots on the right), we observed a clear trend indicating that participants with higher familiarity with the topic were less likely to change their beliefs. However, there is a less obvious trend shown in confirmation bias tendency. We conducted a generalized linear mixed-effects model analysis to quantify the statistical effects (model details were presented in Appendix E.3). The overall result demonstrates that there is a statistically significant difference across the various levels of topic familiarity and confirmation bias tendency on whether participants changed belief or remained unchanged, while the difference is not shown in interface conditions. We briefly describe the mixed effects for each variable in detail as follows.
@@
[…]
Percentage stacked belief changes bar charts between baseline vs. experiment across topic familiarity and confirmation biases tendency
@@
Confirmation bias tendency Confirmation bias tendency shows a statistically significant positive effect on belief change with an estimated coefficient of $0.4338$ ($SE=0.1873$, $z=2.315$ , $p<0.05$). This means that participants with higher levels of confirmation bias tendency are more likely to change their beliefs, which is counterintuitive to our understanding. In a closer examination of belief change patterns across confirmation bias groups, we found a non-linear relationship. For example, the low tendency group demonstrated the highest individual percentage of belief change (25.6%), followed by a decrease in the medium tendency group (19.1%) and then a slight increase in the high tendency group (20.6%). Although the low tendency group showed the highest individual percentage, the combined effect across medium and high groups (39.7% total) exceeded that of the low group, which explains the positive coefficient in our linear model.
Correlation among three independent variables: As parts of the generalized fixed-effect results, there is no moderate correlation among group interface, topic familiarity, and confirmation bias tendency, reflecting that these variables operated relatively independently in influencing belief change.
@@ 5.3.2. Effects of directional belief changes across systems and human factors
We examined how system conditions influence participant directional belief changes, weighted by their confidence level, and explored this variance on topic familiarity and confirmation bias tendency.
@@
Table 4. Directional belief change between baseline and experiment condition: a larger number of participants became less supportive after using the experiment system compared to the baseline. The weighted change for the experiment was significantly lower than that of the baseline, suggesting that the experiment is more likely to lead participants to become less supportive of the stance. The significance levels are represented as p-value, 0 ¡ ‘***’ ¡ 0.001 ¡ ‘**’ ¡ 0.01 ¡ ‘*’ ¡ 0.05
Figure 7. Directional belief changes box plots for the baseline and experimental interfaces across topic familiarity and confirmation biases tendency: Over both factors, while many participants did not alter their beliefs, which reduced the overall magnitude of changes in the box plots, for participants who did change their beliefs, the multi-persona debate system was more likely to lead participants to become less supportive than the baseline.
[…]
Directional belief changes box plots for the baseline and experimental interfaces across topic familiarity and confirmation biases tendency
System comparison: By mapping directional belief changes — less supportive, more supportive, and unchanged — on both systems (Table 4), we found that although many participants did not change their beliefs ($Retrieval_{unchanged}$ = $68.2\%$ , $Debate_{unchanged}$ = $62.5\%$), more participants became less supportive after using the debate than the baseline ($Debate_{less}$ = $24.4\%$ $Retrieval_{less}$ = $11.8\%$). Additionally, by conducting Wilcoxon sign-ranked test on the weighted belief change (in Table 4, the last row, where the mean value ¡0 indicates becoming less supportive and ¿0 indicates becoming more supportive), the weighted belief change for the debate was significantly lower than that of the baseline ($Debate_{weightedchange}$ = $-0.34$ $Retrieval_{weightedchange}$ = $0.30$, $X^{2}=1040.5$, $p<0.05$).
Topic familiarity: By visualizing participant weighted score for belief change across different levels of topic familiarity (Figure 7, the two plots on the left), we found many participants did not alter their beliefs, which reduced the overall magnitude of changes shown in the box plots. However, among participants who did change their beliefs, there is a lowered central tendency (median and the interquartile range) in the experiment compared to the baseline. This expands on earlier findings of significant differences across levels of topic familiarity (Section 5.3.1), demonstrating that, for participants who finally changed their belief, the debate system led them to become less supportive across different familiarity levels. Additionally, we also found the weighted belief change varies depending on different topics (additional analysis and findings are presented in Appendix E.4).
@@ 6. Discussion
6.1. Summary of Key Findings
In this section, we synthesize key findings from Section 5 that address each research question, drawing on both quantitative evidence and qualitative insights. We also contextualize these findings within the existing literature, examining how they align with or diverge from previous work on technology-facilitated critical thinking (as described in Section 2.3) and approaches to mitigating confirmation bias (Section 2.2).
6.1.1. Quantitative insights
Regarding participant experience with the debate system (RQ1), we identified three distinct behavioral patterns through cluster analysis of user interactions (Section 5.1.1). Our approach differs from Zhang et al. (2024), who also implemented LLM multi-persona debates but categorized user behaviors based solely on whether users modified personas or accepted system defaults. Their analysis results in three user groups: interest-driven, system-driven, or mixed. Our classification incorporates a more comprehensive set of interaction metrics and more importantly correlates these behavioral clusters with participant visual attention (Section 5.2.2). By capturing differences in how users engage with diverse perspectives, personalize personas, and manage their reading attention, this approach supports a more nuanced understanding of user behavior using multi-persona debate and its role in reducing confirmation bias.
Regarding which systems are more effective in reducing user confirmation bias (RQ2), our statistical analysis (described in Section 5.2.1) reveals that participants paid more attention to attitude-inconsistent content when using the debate system compared to the baseline system. Regarding the two additional independent variables — topic familiarity and confirmation bias tendency — we found that while topic familiarity did not affect participants’ visual attention on either system, confirmation bias tendency significantly influenced visual attention in the baseline system. These findings suggest that the debate system effectively reduces the Motivational aspect of confirmation bias — increased engagement with attitude-inconsistent content, whereas the baseline system activates the Cognitive aspect of confirmation bias — confirmation bias tendencies.
In terms of the system influence on users’ final belief changes (RQ3), our statistical analysis (Section 5.3) indicates that system conditions did not produce a statistically significant effect on whether participants changed their beliefs (non-directional change). However, there was a notable difference in the direction of changes — more or less supportive to the initial stance on the topic. For example, the debate system was more likely to decrease support, while the retrieval system tended to reinforce supportive beliefs. Further analysis revealed that these changes varied by topic (see Appendix E.4). The retrieval system increased participant support for ‘Animal Testing’ and ‘Space Colonization,’ whereas the debate system decreased support for ‘Federal Corporate Income Tax Raise’ and ‘Violent Video Game Effect on Youth Violence.’ Additionally, there was an overall trend of increased support for ‘Universal Basic Income.’ Moreover, we found that if users were more familiar with the topic, it is less likely for them to change belief.
Reflecting user qualitative responses (Section 5.1.2 and 5.1.3), we assume that the effectiveness of each system in shifting beliefs depends heavily on the nature of the topic and user pre-conception. Compared to prior work, few studies have examined the role of topic differences and explicitly measured user familiarity when evaluating system effectiveness of fostering critical thinking or reducing confirmation bias. Most recent work also relies on one or two topics (Kim et al., 2021; Kriplean et al., 2012; Faridani et al., 2010). By incorporating multiple topics and accounting for user familiarity, our study offers a more nuanced understanding of how passive information presentation (i.e., retrieved content snippets) versus active argumentative formats (i.e., multi-persona debate arguments) interact with users’ pre-existing beliefs. Our study not only highlights the variability in belief shifts across topics but also underscores the need for future research to account for topic-specific effects when designing and evaluating similar systems.
6.1.2. Qualitative insights
Participant qualitative responses complement the behavioral patterns found in the quantitative data, further demonstrating that multi-persona debate can effectively serve as technological interventions to promote critical thinking. Compared to prior work, we identify similar qualitative factors but with nuanced differences.
As highlighted in previous studies, identity plays a crucial role in determining user trust toward diverse opinions. For example, research utilizing crowdsourced opinions found that participants were concerned about the factuality, authority, and reliability of information sources (Kriplean et al., 2012; Kim et al., 2019). Similar concerns arise in our participant feedback regarding the expertise and individuality of LLM personas (Section 5.1.2), as well as the logical robustness and factuality of generated debates (Section 5.1.3).
Additionally, prior work suggests that identity could create a trust paradox (Kim et al., 2021). For example, when users encounter opinions, they often rely on demographic cues to assess an author’s relevant experience and personal stake, which can enhance the perceived credibility of the information. However, this evaluative process also risks triggering representativeness heuristic, where users may inaccurately associate certain demographic characteristics with stereotypical viewpoints.
LLM multi-persona debates carry a risk of reinforcing stereotypes, but this can be partially addressed through carefully crafted persona descriptions that offer depth and context without referencing real individuals. The debate system enables participants to customize persona traits through iterative prompting. This editing process not only allows users to correct potential misalignment between their real-world perspectives and the arguments generated by the LLM, but also encourages a more reflective evaluation by slowing down the judgment process — helping to counteract the influence of the representativeness heuristic. While this cannot fully eliminate inaccurate associations between stereotypical viewpoints and demographic characteristics, it helps mitigate the compounding effects of the representativeness heuristic and confirmation bias, which often reinforce each other in decision-making (Newkirk, 2014).
In summary, the multi-persona debate system significantly reduces confirmation bias by encouraging users to engage more deeply with content that challenges their initial beliefs, as supported by eye-tracking data and behavioral measures. In the following sections, we present additional findings and reflections that, while beyond the primary scope of this study, are closely aligned with directions for future research.